# Safe Multi-Robotic Arm Interaction via 3D Convex Shapes

### v2

@@ 5.2 Centralized-decentralized HOCBF filters
The centralized HOCBF filter employs a single shared filter to process all robots’ nominal input signals collectively, whereas the decentralized filter applies individual safety filters to each robot in the environment. The notion of multi-robot safety in this paper is formally defined in definition 3. As our focus is on inter-robotic arm collision avoidance, this definition does not account for self-collisions. However, analogous HOCBFs for self-collision avoidance can also be formulated by using single-robot dynamics in place of the pairwise dynamics in the multi-robot HOCBF construction described in Sec. 5.1.
@@
In definition 3, the set $\bigcap_{\begin{subarray}{c}k\in\mathcal{P}_{i},l\in\mathcal{P}_{j}\end{subarray}}C^{i_{k}j_{l}}_{1}$ represents the states corresponding to robot configurations in which the collision bodies do not collide. To prevent inter-robot collisions, we require the robots’ states to remain within these safe configurations. Since the relative degree of our inter-robot safety HOCBFs with respect to the pairwise robot dynamics is two, the higher-order sets $C^{i_{k}j_{l}}_{2}$ are necessary to ensure the forward invariance of the sets $C^{i_{k}j_{l}}_{1}$. Specifically, the sets $C^{i_{k}j_{l}}_{2}$ impose velocity-related conditions, guaranteeing that the safety constraints defined by $C^{i_{k}j_{l}}_{1}$ remain dynamically enforceable under appropriate control inputs.
@@
As with centralized-HOCBF-QP, Theorem 1 guarantees that if the multi-robot system starts in a configuration where condition (25) holds and (26) is feasible for all $i\in\mathcal{I}$ at all times, then decentralized-HOCBF-QP ensures safety for all $t\geq t_{0}$. However, the solution to decentralized-HOCBF-QP is suboptimal compared to the centralized-HOCBF-QP solution due to the hard task priorities imposed through the parameter $c_{ij}$ in (26). This also reduces the feasibility set of the control inputs, which may render (26) infeasible for some $i\in\mathcal{I}$. To mitigate this issue, a relaxation method was introduced in decentralized QPs in prior work [4] for first-order CBFs. Following this approach, we propose decentralized-HOCBF-relaxed-QP and establish its theoretical foundation for HOCBFs in Theorem 3.
@@
Let $h$ be a time-invariant HOCBF, and consider $\psi_{k-1}$ with the corresponding safety sets $\mathcal{C}_{k}$ for $k\in\{1,\ldots,r\}$ as defined in (2). Suppose that the class $\mathcal{K}$ functions $\Gamma_{k}$s used to define $\psi_{k}$s are locally Lipschitz. Furthermore, let $\phi(t)$ be continuous and satisfy $\phi(t)\geq 1$ for all $t\geq t_{0}$. Then, any Lipschitz continuous controller $\mathbf{u}:\mathbb{R}^{n}\to\mathbb{R}^{q}$ such that $\mathbf{u}(\mathbf{x})\in K_{\text{RHOCBF}}(\mathbf{x}):=\{\mathbf{u}\in\mathcal{U}:L_{f}\psi_{r-1}(\mathbf{x})+L_{g}\psi_{r-1}(\mathbf{x})\mathbf{u}+\phi(t)\Gamma_{r}(\psi_{r-1}(\mathbf{x}))\geq 0\}$ will render the set $\mathcal{C}:=\bigcap_{k=1}^{r}\mathcal{C}_{k}$ forward invariant for the system (1).
@@
The solution is unique because $\Gamma_{r}$ is locally Lipschitz and $\phi$ is continuous. Since $\mathbf{u}(\mathbf{x})\in K_{\text{RHOCBF}}(\mathbf{x})$, we have $\dot{\psi}_{r-1}(\mathbf{x})=L_{f}\psi_{r-1}(\mathbf{x})+L_{g}\psi_{r-1}(\mathbf{x})\mathbf{u}\geq-\phi(t)\Gamma_{r}(\psi_{r-1}(\mathbf{x}))$. From (27) and Comparison Lemma [25], it follows that
@@
Assuming $\mathbf{x}(0)\in\mathcal{C}$, the definition of $\mathcal{C}$ implies $\psi_{k}(\mathbf{x}(0))\geq 0,\,0\leq k\leq r-1$. Since $\phi(t)\geq 1$, $\int_{0}^{t}\phi(\tau)d\tau$ is monotonically increasing with $t$ and $\int_{0}^{t}\phi(\tau)d\tau\rightarrow\infty$ as $t\to\infty$. By the properties of $\mathcal{K}\mathcal{L}$ functions, for a fixed $u\geq 0$, $\sigma(u,v)\rightarrow 0$ as $v\rightarrow\infty$ and as it is decreasing, $\psi_{r-1}(\mathbf{x})\geq 0$ for all $t\geq 0$. Given the recursive definition $\psi_{k}(\mathbf{x})=\left(\frac{d}{dt}+\Gamma_{k}\right)\psi_{k-1}(\mathbf{x})$, if $\psi_{k}(\mathbf{x})\geq 0$, then $\dot{\psi}_{k-1}(\mathbf{x})\geq-\Gamma_{k}({\psi}_{k-1}(\mathbf{x}))$. Since $\Gamma_{k}$ is a class $\mathcal{K}$ function and $\mathbf{x}(0)\in\mathcal{C}$, it follows by induction that $\psi_{k}(\mathbf{x})\geq 0$ for $0\leq k\leq r-1$. This establishes that $K_{\text{RHOCBF}}(\mathbf{x})$ renders the set forward-invariant for system (1). ∎
As the only constraint on $\phi(t)$ is that it must be continuous, we follow [4] and introduce it as an optimization parameter in the decentralized QPs. The optimization problem includes the hyperparameter $\lambda$, which controls the trade-off between increasing $\phi(t)$ and adhering to the original objective, and is formulated for robot $i\in\mathcal{I}$ as follows:
@@
In the centralized filter, QP infeasibility, i.e. an empty feasible set, can theoretically arise, although we did not encounter this situation in any experiment. A typical place where it may become infeasible is in highly cluttered environments where a robot’s movement is prevented by multiple surrounding robots. Even when the centralized QP remains feasible, one or more QPs in the decentralized filters may become infeasible under the same scenario. This occurs because the decentralized formulation shrinks the feasible sets to decouple the shared problem into independent subproblems. When QP infeasibility is observed, a backup controller that halts all robots by applying accelerations opposite to their joint velocities is activated. Since the safety filters already bound joint speeds, this braking action stops the robots before contact. In practice, we observed no QP infeasibility with the centralized filter and only a few cases with the decentralized filters; see Sec. 7 for details.
@@ 6 Computational challenges in computing HOCBFs for multi-robotic arm safety
In this section, we address the computational challenges of employing the HOCBFs in multi-robot environments. During the experiments presented in Sec. 7, we observed that the primary computational overhead in HOCBF computations arises from computing $L_{f}\psi_{1}^{i_{k}ij}(\mathbf{x}_{ij})$. While this overhead does not hinder computation of HOCBFs that prevent a single robot from colliding with static or dynamic obstacles as in [3], it poses significant computational challenges in multi-robot environments due to the large number of HOCBFs required to ensure multi-robot safety. For instance, while 14 HOCBFs are sufficient to protect the 7 links of a robotic arm from two obstacles, 48 HOCBFs are required to protect the last 4 links of the three robotic arms from colliding with each other. Furthermore, the need to define $\boldsymbol{\theta}$ as a 14-dimensional vector for two moving objects—compared to 7 dimensions for a moving robot and a static obstacle—further amplifies the computational burden of HOCBF computations.
@@
$L_{f}\psi_{1}^{i_{k}ij}(\mathbf{x}_{ij})$ is computed as in (17). The primary bottleneck in this computation stems from calculating the Hessian $\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\in\mathbb{R}^{14\times 14}$. This Hessian is required to compute the term $\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}^{T}\left(\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}$. We refer to this term as Hessian contribution throughout the remainder of the paper. To alleviate the computational burden, we propose approximating $\left(\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}=\frac{d}{dt}\left(\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)$ via numerical differentiation. This numerical derivative is obtained by using the gradient computed at the current time step together with gradients from previous time steps. The approximated Hessian contribution is then obtained by taking the dot product of this numerical derivative with $\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}$. Note that the gradient $\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\dot{\boldsymbol{\theta}}_{i_{k}j_{l}})$ is already computed as part of other terms in the HOCBF formulation, so no additional gradient evaluation is necessary. Since real robot sensor readings can be noisy and susceptible to environmental perturbations, we use a robust numerical differentiation filter, the Savitzky-Golay filter [26], in our methodology, where its parameters can be tuned according to the noise level.
@@ 7 Experiments
We designed both simulation and real-world experiments to evaluate the proposed method. In real-world experiments, we demonstrated the applicability of the methodology under real-world conditions and showed the effect of utilizing numerical derivation in Hessian contribution on the control update rate, while in simulations, we showcased its scalability to scenarios involving more than two robots and various path-following tasks. Additionally, we exhaustively compared centralized, decentralized, and relaxed-decentralized HOCBF filters in the simulations. In all experiments, feedback linearization was used, and the safety filters were formulated in terms of joint accelerations. An identity matrix was employed for the matrices $\hat{\textbf{P}}_{i}$ and $\hat{\textbf{P}}$ in (32) and (33). The parameter $c_{ij}$ in the decentralized and relaxed-decentralized filters was set to $0.5$. $\lambda$ in the relaxed-decentralized filter was chosen as $10$. We utilized OSQP [27] to solve the QP problems. For the nominal controller, we incorporated an inverse dynamics controller as the primary objective for trajectory-following in the operational space.
In the experiments, we utilized linear class $\mathcal{K}$ functions for multi-robot safety HOCBFs, i.e $\Gamma_{w}(\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij}))=\Gamma_{w}\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij})$ in (11). We set $\Gamma_{1}=15$ and $\Gamma_{2}=15$ for these functions and selected $\alpha_{0}=1.03$ as the parameter in (11). Similarly, we used linear class $\mathcal{K}$ functions for the joint angle and velocity HOCBFs, setting $\Gamma_{1}^{\underline{\mathbf{q}}_{i}}=\Gamma_{2}^{\underline{\mathbf{q}}_{i}}=\Gamma_{1}^{\overline{\mathbf{q}}_{i}}=\Gamma_{2}^{\overline{\mathbf{q}}_{i}}=20$ and $\Gamma_{1}^{\underline{\dot{\mathbf{q}}_{i}}}=\Gamma_{1}^{\overline{\dot{\mathbf{q}}}_{i}}=10$.

### v3

@@ 5.2 Centralized-decentralized HOCBF filters
The centralized HOCBF filter employs a single shared filter to process all robots’ nominal input signals collectively, whereas the decentralized filter applies individual safety filters to each robot in the environment. The notion of multi-robot safety in this paper is formally defined in Definition 3. As our focus is on inter-robotic arm collision avoidance, this definition does not account for self-collisions. However, analogous HOCBFs for self-collision avoidance can also be formulated by using single-robot dynamics in place of the pairwise dynamics in the multi-robot HOCBF construction described in Sec. 5.1. Similarly, the proposed framework can be extended to dynamic obstacles in the environment by treating each obstacle as an additional agent. Pairwise HOCBFs can then be constructed between the obstacle components and the robot links and end-effectors, and the resulting constraints can be integrated into the safety filters.
@@
In Definition 3, each set $C^{i_{k}j_{l}}_{1}$ is defined as the superlevel set of $\psi^{i_{k}j_{l}}_{0}(\mathbf{x}_{ij})=\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})-\alpha_{0}$, where nonnegative values indicate that the corresponding robot link pair $(i_{k},j_{l})$ remains safely separated. Definition 3 enforces this condition for all link and end-effector pairs belonging to different robotic arms. Therefore, in terms of collision avoidance, it defines multi-robotic arm safety. Since the relative degree of our inter-robot safety HOCBFs with respect to the pairwise robot dynamics is two, the higher-order sets $C^{i_{k}j_{l}}_{2}$ are necessary to ensure the forward invariance of the sets $C^{i_{k}j_{l}}_{1}$. Specifically, the sets $C^{i_{k}j_{l}}_{2}$ impose velocity-related conditions, guaranteeing that the safety constraints defined by $C^{i_{k}j_{l}}_{1}$ remain dynamically enforceable under appropriate control inputs.
@@
The relationship between the centralized and decentralized filters can be explained as follows. Assuming that $\mathbf{P}$ in centralized-HOCBF-QP has a block-diagonal structure, where each block $\mathbf{P}_{i}$ corresponds to robot $i$, the cost function in centralized-HOCBF-QP can be written as $\sum_{i\in\mathcal{I}}(\mathbf{u}_{i}-\mathbf{u}_{\mathrm{nom}_{i}})^{T}\mathbf{P}_{i}(\mathbf{u}_{i}-\mathbf{u}_{\mathrm{nom}_{i}})$. To separate the optimization constraints for each robot, the multi-robot safety HOCBF constraints in centralized-HOCBF-QP are decoupled through the introduction of responsibility coefficients $c_{ij}$. This decomposition, together with the independence of the cost function terms across robots, isolates each robot’s control input in the optimization problem from those of the other robots. As a result, each robot can solve its own QP locally and in parallel, referred to as decentralized-HOCBF-QP, enforcing its share of the safety constraints while considering only its control input.
As with centralized-HOCBF-QP, Theorem 1 guarantees that if the multi-robot system starts in a configuration where condition (25) holds and (26) is feasible for all $i\in\mathcal{I}$ at all times, then decentralized-HOCBF-QP ensures safety for all $t\geq t_{0}$. However, the solution to decentralized-HOCBF-QP is suboptimal compared to the centralized-HOCBF-QP solution due to the decoupling of the shared multi-robot safety constraints through the responsibility coefficients $c_{ij}$ in (26). This also reduces the feasibility set of the control inputs, which may render (26) infeasible for some $i\in\mathcal{I}$. To mitigate this issue, a relaxation method was introduced in decentralized QPs in prior work wang2017safety for first-order CBFs. Following this approach, we propose decentralized-HOCBF-relaxed-QP and establish its theoretical foundation for HOCBFs in Theorem 3.
@@
Let $h$ be a time-invariant HOCBF, and consider $\psi_{k-1}$ with the corresponding safety sets $C_{k}$ for $k\in\{1,\ldots,r\}$ as defined in (2). Suppose that the class $\mathcal{K}$ functions $\Gamma_{k}$s used to define $\psi_{k}$s are locally Lipschitz. Furthermore, let $\phi(t)$ be continuous and satisfy $\phi(t)\geq 1$ for all $t\geq t_{0}$. Then, any Lipschitz continuous controller $\mathbf{u}:\mathbb{R}^{n}\to\mathbb{R}^{q}$ such that $\mathbf{u}(\mathbf{x})\in K_{\text{RHOCBF}}(\mathbf{x}):=\{\mathbf{u}\in\mathcal{U}:L_{f}\psi_{r-1}(\mathbf{x})+L_{g}\psi_{r-1}(\mathbf{x})\mathbf{u}+\phi(t)\Gamma_{r}(\psi_{r-1}(\mathbf{x}))\geq 0\}$ will render the set $\mathcal{C}:=\bigcap_{k=1}^{r}C_{k}$ forward invariant for the system (1).
@@
The solution is unique because $\Gamma_{r}$ is locally Lipschitz and $\phi$ is continuous. Since $\mathbf{u}(\mathbf{x})\in K_{\text{RHOCBF}}(\mathbf{x})$, we have $\dot{\psi}_{r-1}(\mathbf{x})=L_{f}\psi_{r-1}(\mathbf{x})+L_{g}\psi_{r-1}(\mathbf{x})\mathbf{u}\geq-\phi(t)\Gamma_{r}(\psi_{r-1}(\mathbf{x}))$. From (27) and Comparison Lemma khalil1996nonlinear , it follows that
@@
Assuming $\mathbf{x}(0)\in\mathcal{C}$, the definition of $\mathcal{C}$ implies $\psi_{k}(\mathbf{x}(0))\geq 0,\,0\leq k\leq r-1$. Since $\phi(t)\geq 1$, $\int_{0}^{t}\phi(\tau)d\tau$ is monotonically increasing with $t$ and $\int_{0}^{t}\phi(\tau)d\tau\rightarrow\infty$ as $t\to\infty$. By the properties of $\mathcal{K}\mathcal{L}$ functions, for a fixed $u\geq 0$, $\sigma(u,v)\rightarrow 0$ as $v\rightarrow\infty$ and as it is decreasing, $\psi_{r-1}(\mathbf{x})\geq 0$ for all $t\geq 0$. Given the recursive definition $\psi_{k}(\mathbf{x})=\dot{\psi}_{k-1}(\mathbf{x})+\Gamma_{k}\left(\psi_{k-1}(\mathbf{x})\right)$, if $\psi_{k}(\mathbf{x})\geq 0$, then $\dot{\psi}_{k-1}(\mathbf{x})\geq-\Gamma_{k}({\psi}_{k-1}(\mathbf{x}))$. Since $\Gamma_{k}$ is a class $\mathcal{K}$ function and $\mathbf{x}(0)\in\mathcal{C}$, it follows by induction that $\psi_{k}(\mathbf{x})\geq 0$ for $0\leq k\leq r-1$. This establishes that $K_{\text{RHOCBF}}(\mathbf{x})$ renders the set forward-invariant for system (1). ∎
As the only constraint on $\phi(t)$ is that it must be continuous, we follow wang2017safety and introduce it as an optimization parameter in the decentralized QPs. The optimization problem includes the hyperparameter $\lambda$, which controls the trade-off between increasing $\phi(t)$ and adhering to the original objective, and is formulated for robot $i\in\mathcal{I}$ as follows:
@@
In the centralized filter, QP infeasibility, i.e. an empty feasible set, can theoretically arise, although we did not encounter this situation in any experiment. A typical place where it may become infeasible is in highly cluttered environments where a robot’s movement is prevented by multiple surrounding robots. Even when the centralized QP remains feasible, one or more QPs in the decentralized filters may become infeasible under the same scenario. This occurs because the decentralized formulation shrinks the feasible sets to decouple the shared problem into independent subproblems. When QP infeasibility is observed, a fail-safe mode that halts all robots by applying accelerations opposite to their joint velocities is activated. Since the safety filters already bound joint speeds, this braking action stops the robots before contact. However, activating the fail-safe mode interrupts the high-level task being executed by the nominal controller and requires new high-level planning before resuming operation. In practice, we observed no QP infeasibility with the centralized filter and only a few cases with the decentralized filters; see Sec. 7 for details.
@@ 6 Computational challenges in computing HOCBFs for multi-robotic arm safety
In this section, we address the computational challenges of employing the HOCBFs in multi-robot environments. During the experiments presented in Sec. 7, we observed that the primary computational overhead in HOCBF computations arises from computing $L_{f}\psi_{1}^{i_{k}j_{l}}(\mathbf{x}_{ij})$. While this overhead does not hinder computation of HOCBFs that prevent a single robot from colliding with static or dynamic obstacles as in wei2024collision , it poses significant computational challenges in multi-robot environments due to the large number of HOCBFs required to ensure multi-robot safety. For instance, while 14 HOCBFs are sufficient to protect the 7 links of a robotic arm from two obstacles, 48 HOCBFs are required to protect the last 4 links of the three robotic arms from colliding with each other. Furthermore, the need to define $\boldsymbol{\theta}$ as a 14-dimensional vector for two moving objects—compared to 7 dimensions for a moving robot and a static obstacle—further amplifies the computational burden of HOCBF computations.
@@
$L_{f}\psi_{1}^{i_{k}j_{l}}(\mathbf{x}_{ij})$ is computed as in (17). The primary bottleneck in this computation stems from calculating the Hessian $\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\in\mathbb{R}^{14\times 14}$. This Hessian is required to compute the term $\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}^{T}\left(\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}$. We refer to this term as Hessian contribution throughout the remainder of the paper. To alleviate the computational burden, we propose approximating $\left(\nabla^{2}_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}=\frac{d}{dt}\left(\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)$ via numerical differentiation. This numerical derivative is obtained by using the gradient computed at the current time step together with gradients from previous time steps. The approximated Hessian contribution is then obtained by taking the dot product of this numerical derivative with $\dot{\boldsymbol{\theta}}_{i_{k}j_{l}}$. Note that the gradient $\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})$ is already computed as part of other terms in the HOCBF formulation, so no additional gradient evaluation is necessary. Since real robot sensor readings can be noisy and susceptible to environmental perturbations, we use a robust numerical differentiation filter, the Savitzky-Golay filter savitzky1964smoothing , in our methodology, where its parameters can be tuned according to the noise level.
The Savitzky–Golay filter is characterized by two hyperparameters: the window length and the polynomial order. It fits a polynomial of the chosen order to the most recent window of data using least squares and computes the derivative from the fitted curve. These hyperparameters affect the accuracy of the estimated Hessian contribution. An excessively large window length produces overly smoothed and lagged derivatives that fail to capture rapid changes. Conversely, a small window makes the filter sensitive to the noise in the signal. Higher-order polynomials capture complex trends but are more noise-sensitive, while lower orders offer better robustness against noise. In our implementation, we use a second-degree polynomial and select the window length based on the expected noise level. Our proposed safety method operates robustly for a reasonable range of these hyperparameter values, indicating low sensitivity to their precise selection. Nonetheless, inappropriate hyperparameter choices can amplify derivative estimation errors, distort the estimated Hessian contribution, and ultimately compromise safety.
@@ 7 Experiments
We designed both simulation and real-world experiments to evaluate the proposed method. In real-world experiments, we demonstrated the applicability of the methodology under real-world conditions and showed the effect of utilizing numerical derivation in Hessian contribution on the control update rate, while in simulations, we showcased its scalability to scenarios involving more than two robots and various path-following tasks. Additionally, we exhaustively compared centralized, decentralized, and relaxed-decentralized HOCBF filters in the simulations. In all experiments, feedback linearization was used, and the safety filters were formulated in terms of joint accelerations. An identity matrix was employed for the matrices $\hat{\textbf{P}}_{i}$ and $\hat{\textbf{P}}$ in (32) and (33). The parameter $c_{ij}$ in the decentralized and relaxed-decentralized filters was set to $0.5$. $\lambda$ in the relaxed-decentralized filter was chosen as $10$. We utilized OSQP osqp to solve the QP problems. For the nominal controller, we incorporated an inverse dynamics controller as the primary objective for trajectory-following in the operational space.
In the experiments, we utilized linear class $\mathcal{K}$ functions for multi-robot safety HOCBFs, i.e $\Gamma_{w}(\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij}))=\Gamma_{w}\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij})$ in (11). We set $\Gamma_{1}=15$ and $\Gamma_{2}=15$ for these functions. Similarly, we used linear class $\mathcal{K}$ functions for the joint angle and velocity HOCBFs, setting $\Gamma_{1}^{\underline{\mathbf{q}}_{i}}=\Gamma_{2}^{\underline{\mathbf{q}}_{i}}=\Gamma_{1}^{\overline{\mathbf{q}}_{i}}=\Gamma_{2}^{\overline{\mathbf{q}}_{i}}=20$ and $\Gamma_{1}^{\underline{\dot{\mathbf{q}}_{i}}}=\Gamma_{1}^{\overline{\dot{\mathbf{q}}}_{i}}=10$. We selected $\alpha_{0}=1.03$ as the parameter in (11). This value provided a small but sufficient safety margin to prevent collisions while avoiding excessive conservatism in the experiments.